# Window-based Channel Attention for Wavelet-enhanced Learned Image Compression

### v2

@@ 1 Introduction
Convolutional neural networks (CNN) have demonstrated competence in LIC. Ballé et al. [2] built a basic framework for CNN-based LIC, which was further extended by incorporating VAE and hyper-prior modules to enhance the performance [3]. Cheng et al. [10] proposed auto-regressive models [34] and Gaussian mixture models (GMM) to improve the entropy coding performance. Transformer-based models, leveraging the advantages of Transformer architectures in computer vision tasks, have emerged as promising alternatives in LIC [55, 40, 56, 24, 27]. These models, such as Swin-Transformer-based model [55] and parallel bidirectional context Transformer-based model [40], have demonstrated superior compression performance over classic CNN-based approaches.
@@
Other works introduced Discrete Wavelet Transform (DWT) to effectively enlarge receptive fields for computer vision tasks [29, 28, 48, 21, 54, 27]. DWT decomposes the image into sub-images with different frequency properties, functioning as a parameter-free down-sampling method to enlarge receptive fields. DWT has also been explored in [32, 33, 41] to efficiently encode images and optimize bit allocation for image details in LIC. In this work, we propose to use DWT to further enlarge receptive fields for our window-based channel attention as illustrated in Figs.1(c) and 1(d), thereby improving performances of our LIC system.
@@ 2.1.1 CNN-based Models
In recent years, there has been a rapid advancement in LIC [2, 3, 34, 36, 35, 10, 37, 18, 53], marked by notable breakthroughs. Ballé et al. [2] built a basic CNN-based framework for end-to-end LIC models. Subsequently, they extended their work by introducing a VAE architecture and incorporating a hyper-prior module to enhance the capabilities of LIC [3]. In pursuit of improved entropy modeling, an auto-regressive component was incorporated in [34], drawing inspiration from the success of auto-regressive priors. Cheng et al. [10] substituted the Single Gaussian Model (SGM) in the entropy model with a Gaussian Mixture Model (GMM), which involved the integration of residual blocks and a simplified attention module. Addressing the computational overhead of the context model, He et al. [18] introduced a checkerboard context model designed for parallel computing. Minnen et al. [35] reduced computational costs by employing a channel-wise context. Beyond advancements in entropy modeling, researchers have explored diverse CNN architectures to elevate feature extraction for image compression. Liu et al. [26] introduced non-local residual blocks to capture local and global correlations, leveraging attention masks to allocate bits intelligently based on feature importance. Xie et al. [53] leveraged invertible neural networks (INNs) to enhance overall performance. Pan et al. [37] introduced the content adaptive channel dropping (CACD), aiming to improve the content adaptability on both latents and the decoder.
@@ 3.1 Problem Formulation
Given an input image $x$, the analysis transform $g_{a}$ maps it to a latent representation $y$, which is then quantized into $\lceil y\rfloor$ by $Q$, and we employ a range coder to encode it losslessly. According to [34], we encode $\lceil y-\mu\rfloor$ instead of $\lceil y\rfloor$, where $\mu$ is the mean estimated by the entropy model. Subsequently, we reconstruct the coded $\hat{y}$ as $\lceil y-\mu\rfloor+\mu$, and it is mapped back to the reconstructed image $\hat{x}$ using the synthesis transform $g_{s}$. The main pipeline is formulated as:
$y=g_{a}(x;\phi),\quad\hat{y}$ $=Q(y-\mu)+\mu,\quad\hat{x}=g_{s}(\hat{y};\theta),$ (1)
@@
Image compression considers both rate and distortion, so the loss function of LIC is defined as:
@@ 3.2 Method Overview
Figure 2: The overall architecture of our model. SCH is Space-Channel Hybrid block. DWT is the discrete wavelet transform and IDWT is the inverse transform. $\downarrow$ means down-sampling and $\uparrow$ means up-sampling. RB is Residual Block. RBS is Residual Block with Stride. RBU is Residual Block Up-sampling.

### v3

@@ 1 Introduction
Convolutional neural networks (CNN) have demonstrated competence in LIC. Ballé et al. [2] built a basic framework for CNN-based LIC, which was further extended by incorporating VAE and hyper-prior modules to enhance the performance [3]. To improve the entropy coding performance, auto-regressive models [33] and Gaussian mixture models (GMM) [10] have been proposed. Transformer-based models, leveraging the advantages of Transformer architectures in computer vision tasks, have emerged as promising alternatives in LIC [54, 39, 55, 23, 26]. These models, such as Swin-Transformer-based model [54] and parallel bidirectional context Transformer-based model [39], have demonstrated superior compression performance over classic CNN-based approaches.
@@
Other works introduced Discrete Wavelet Transform (DWT) to effectively enlarge receptive fields for computer vision tasks [28, 27, 47, 20, 53, 26]. DWT decomposes the image into sub-images with different frequency properties, functioning as a parameter-free down-sampling method to enlarge receptive fields. DWT has also been explored in [31, 32, 40] to efficiently encode images and optimize bit allocation for image details in LIC. In this work, we propose to use DWT to further enlarge receptive fields for our window-based channel attention as shown in Figs.1(c) and 1(d), thereby improving the performance of our LIC system.
@@ 2.1.1 CNN-based Models
In recent years, there have been notable breakthroughs in CNN-based LIC [2, 3, 33, 35, 34, 10, 36, 17, 52]. Ballé et al. [2] built a basic CNN-based framework for end-to-end LIC models. They extended the work by introducing a VAE architecture and incorporating a hyper-prior module to enhance the capabilities of LIC [3]. Drawing inspiration from the success of auto-regressive priors, Minnen et al. [33] introduced an auto-regressive component to improve entropy modeling. Cheng et al. [10] substituted the Single Gaussian Model (SGM) in the entropy model with a Gaussian Mixture Model (GMM), which involved the integration of residual blocks and a simplified attention module. Addressing the computational overhead of the context model, He et al. [17] introduced a checkerboard context model designed for parallel computing. Minnen et al. [34] reduced computational costs by employing a channel-wise context. Beyond advancements in entropy modeling, researchers have explored diverse CNN architectures to elevate feature extraction for LIC. Liu et al. [25] introduced non-local residual blocks to capture local and global correlations, leveraging attention masks to allocate bits intelligently based on feature importance. Xie et al. [52] exploited invertible neural networks (INNs) to enhance overall performance. Pan et al. [36] introduced the content adaptive channel dropping (CACD), aiming to improve the content adaptability on both latents and the decoder.
@@ 3.1 Problem Formulation
Given an input image $x$, the analysis transform $g_{a}$ maps it to a latent representation $y$, which is then quantized into $\lceil y\rfloor$ by $Q$, and we employ a range coder to encode it losslessly. According to [33], we encode $\lceil y-\mu\rfloor$ instead of $\lceil y\rfloor$, where $\mu$ is the mean estimated by the entropy model. Suppose $y_{q}=\lceil y-\mu\rfloor$, we then reconstruct the coded $\hat{y}$ as $y_{q}+\mu$, and it is mapped back to the reconstructed image $\hat{x}$ using the synthesis transform $g_{s}$. The main pipeline is formulated as:
$y=g_{a}(x;\phi),\quad y_{q}=Q(y-\mu)$ $,\quad\hat{y}=y_{q}+\mu,\quad\hat{x}=g_{s}(\hat{y};\theta),$ (1)
@@
LIC considers both rate and distortion, so the loss function is defined as:
@@ 3.2 Method Overview
Figure 2: The overall architecture of our model. SCH is Space-Channel Hybrid block. DWT is the discrete wavelet transform and IDWT is the inverse transform. $\downarrow$ means down-sampling and $\uparrow$ means up-sampling. RB is Residual Block. RBS is Residual Block with Stride. RBU is Residual Block Up-sampling. RE is Range Encoder and RD is Range Decoder.